OpenStudy (anonymous):

how to determine the points of this equation? 2x-3y=-6 -3y+2x=3

OpenStudy (***[isuru]***):

Hi, r u trying to mention "intersection points" ?

OpenStudy (anonymous):

plot on coordinate plane

OpenStudy (anonymous):

im trying to find ordered pairs to determine its solution

OpenStudy (anonymous):

ordered pairs of each equation.

OpenStudy (anonymous):

Do you know method of elimination or of substitution?

OpenStudy (***[isuru]***):

@wio the given equations can't be solved ?

OpenStudy (***[isuru]***):

http://fooplot.com/plot/svcjthxk05

OpenStudy (anonymous):

yes.. but i need ordered pairs so i can graph it to determine if it has no solution, ifinie solution or exactly one solution

OpenStudy (***[isuru]***):

ah! u want to find co-ordinates for each functions to plot them ?

OpenStudy (anonymous):

yes

OpenStudy (***[isuru]***):

take one function at a time and substitute values to "x" and get the result ; which is the value of y. i.e - 2x-3y=-6 when x is 1 y= 8/3 like wise... but there r many ways to show that these 2 has no solutions ....

OpenStudy (anonymous):

i need ordered pairs, :(

OpenStudy (***[isuru]***):

here r the ord ers pairs 2x-3y=-6 (-3 , 0) (0 , 2) (3 , 4) (-6 , -2) these r more than enough to plot graph 1 for the second one ... -3y+2x=3 (0 , -1) (-3 , -3) (-6 , -5) (3, 1) (6, 3) hope this will help ya!!

OpenStudy (anonymous):

thanks! :D

OpenStudy (***[isuru]***):

u r welcome!!!!